OpenStudy (anonymous):

Choose the equation of the line passing through the point (2, -4) and parallel to y = 3x - 6. y = 3x - 10 y = 3x - 6 y = 3x - 2 y = 3x - 12

OpenStudy (anonymous):

y=3x-6 so the slope is 3 y-(-4)=3(x-2) y+4=3x-6 y=3x-10

OpenStudy (anonymous):

How did you get that answer

OpenStudy (anonymous):

the coefficient of x is the slope of the equation. u should write it like this equation: y-y'=m(x-x') m is slope and y' and x' are given (2,-4) so u'll have: y-(-4)=3(x-2)

OpenStudy (anonymous):

thanks